Shares in Potash Inc. are currently trading for $14.83 per share. Potash paid a dividend on December 31st (yesterday). Today is January 1st. The dividend next year (on December 31st) is expected to be $0.83. You plan to sell your shares after the dividend next year and expect to receive $15.52 when you sell them. Denote the anticipated return on your Potash Shares as R. Potash’s beta is 0.78, the expected return on the market is 8.8% and T-Bills are expected to earn 4.1%. What is the Jensen’s Alpha coefficient for your investment in Potash Inc? (Jensen’s alpha is just the difference between the anticipated return, R, and the equilibrium expected return, E(R)).

A) 0%

B) *2.5%

C) 6.1%

D) 7.8%

E) 10.3%

Looking for the process work to understand why the solution is B) 2.5%.
Answer #1

Expected return=(Dividend in Year 1+Price in Year 1)/Price now-1=(0.83+15.52)/14.83-1=10.2495%

Required return=risk free rate+beta*(market return-risk free rate)=4.1%+0.78*(8.8%-4.1%)=7.7660%

Alpha=Expected return-required return=10.2495%-7.7660%=2.4835%